### v1

@@ Abstract
Deep neural networks are often seen as different from other model classes by defying conventional notions of generalization. Popular examples of anomalous generalization behaviour include benign overfitting, double descent, and the success of overparametrization. We argue that these phenomena are not distinct to neural networks, or particularly mysterious. Moreover, this generalization behaviour can be intuitively understood, and rigorously characterized using long-standing generalization frameworks such as PAC-Bayes and countable hypothesis bounds. We present soft inductive biases as a key unifying principle in explaining these phenomena: rather than restricting the hypothesis space to avoid overfitting, embrace a flexible hypothesis space, with a soft preference for simpler solutions that are consistent with the data. This principle can be encoded in many model classes, and thus deep learning is not as mysterious or different from other model classes as it might seem. However, we also highlight how deep learning is relatively distinct in other ways, such as its ability for representation learning, phenomena such as mode connectivity, and its relative universality.
@@ 1 Introduction
We will also treat all of these phenomena collectively, through a unifying notion of soft inductive biases. While inductive biases are often thought of as restriction biases — constraining the size of a hypothesis space for improved data efficiency and generalization — there is no need for restriction biases. Instead, we can embrace an arbitrarily flexible hypothesis space, combined with soft biases that express a preference for certain solutions over others, without entirely ruling out any solution, as illustrated in Figure 3. Frameworks such as PAC-Bayes are entirely consistent with this view of inductive biases, capable of producing non-vacuous generalization bounds on models with even billions of parameters, as long as these models have a prior preference for certain solutions over others (Lotfi et al., 2024b). Broadly speaking, a large hypothesis space, combined with a preference for simple solutions, provides a provably useful recipe for good performance, as in Figure 2.
There are also other phenomena of recent interest, such as scaling laws and grokking, which are not our focus, because these are not typically treated as inconsistent with generalization theory, or distinct to neural networks. However, we note the PAC-Bayes and countable hypothesis generalization frameworks of Section 3 are also consistent with LLMs, and even Chinchilla scaling laws (Hoffmann et al., 2022; Finzi et al., 2025). Moreover, deep learning of course is different and mysterious in other ways. In Section 8, we discuss relatively distinctive features of deep neural networks, such as representation learning, mode connectivity, and broadly successful in-context learning.
@@ 2 Soft Inductive Biases
In Residual Pathway Priors (RPP) (Finzi et al., 2021), it was shown that a soft bias for equivariance constraints was often as effective as a model that had been perfectly constrained for a given problem. For example, a soft bias for rotation equivariance would work as well as a rotationally equivariant model for molecules, which are rotation invariant. After exposure to only a very small amount of data, the soft bias would converge to near-perfect rotation equivariance, since the model is encouraged (but not constrained) to represent the data with symmetries, and it can do so exactly, even with a small amount of data. Moreover, in cases where the data only contained an approximate symmetry, or no symmetry at all, the RPP approach would significantly outperform a model with hard symmetry constraints.
@@ 3 Generalization Frameworks
It turns out these phenomena are also formally characterized by generalization frameworks that have existed for many decades, including PAC-Bayes (McAllester, 1999; Guedj, 2019; Alquier et al., 2024) and simple countable hypothesis bounds (Valiant, 1984; Shalev-Shwartz & Ben-David, 2014). We introduce these frameworks in Section 3.1. We then define effective dimensionality in Section 3.2 which we will return to later in the paper for intuition. Finally we introduce frameworks that do not describe these phenomena in Section 3.3, but have greatly impacted the conventional wisdom in thinking about generalization.
@@ 3.1 PAC-Bayes and countable hypothesis bounds
This bound is related to the finite hypothesis bound, but includes a prior $P(h)$ and a countable rather than finite hypothesis space (Ch 7.3, Shalev-Shwartz & Ben-David, 2014). We can think of the prior as a weighting function that weights certain hypotheses more highly than others. Importantly, we can use any prior to evaluate the bound: it need not have generated the true hypothesis for the data, contain the true hypothesis, or even be used by the model that is trained to find some hypothesis $h^{*}$. If the model uses a prior quite different from the prior used to evaluate Eq. (1), then the bound will simply become loose. We including an elementary proof of this bound in Appendix C.
We can derive informative bounds through a Solomonoff prior $P(h)\leq 2^{-K(h|A)}$ (Solomonoff, 1964) where $K$ is the prefix-free Kolmogorov complexity of $h$ taking as input model architecture $A$. Substituting this prior into Eq. (1),
@@ 3.2 Effective Dimensionality
Solutions with lower effective dimensionality are flatter, meaning that the associated parameters can be perturbed without significantly increasing the loss. Flatness is not the only factor influencing generalization, and flatness as measured by the Hessian is not parametrization invariant (like SGD, $\ell_{2}$ regularization, and many standard procedures), meaning it is easy to find and construct examples where flatter solutions do not generalize better (e.g., Dinh et al., 2017). On the other hand, the connection between flatness and generalization is not a spurious empirical association. We have a mechanistic understanding of why flatness can lead to better generalization: flatter solutions are more compressible, have better Occam factors, tend to lead to wider decision boundaries, and tighter generalization bounds (Hinton & Van Camp, 1993; Hochreiter & Schmidhuber, 1997; MacKay, 2003; Keskar et al., 2016; Izmailov et al., 2018; Foret et al., 2020; Maddox et al., 2020).
@@
We will often return to effective dimensionality for intuition when discussing generalization phenomena.
@@ 3.3 Other Generalization Frameworks
Rademacher complexity (Bartlett & Mendelson, 2002) exactly measures the ability for a model to fit uniform $\{+1,-1\}$ random noise. Similarly, the VC dimension (Vapnik et al., 1994) measures the largest integer $d$ such that the hypothesis space $\mathcal{H}$ can fit (“shatter”) any set of $d$ points with $\{+1,-1\}$ labels. The fat-shattering dimension (Alon et al., 1997) $\text{fat}_{\gamma}(\mathcal{H})$ refines the VC dimension to fitting (“shattering”) labels by some margin $\gamma$. Unlike PAC-Bayes, all of these frameworks penalize the size of the overall hypothesis space $\mathcal{H}$, suggesting a prescription for restriction biases, rather than the soft inductive biases of Section 2. We discuss these frameworks further in Appendix B, with a comparative summary in Table 1.
@@ 4 Benign Overfitting
There is by now a large body of work studying and reproducing benign overfitting with other model classes. Yet the conventional wisdom of benign overfitting as a mysterious and deep learning specific phenomenon, one that still requires rethinking generalization, persists. It is not our intention, nor would it be possible, to cover all of this work here, but we note some of the key developments. Dziugaite & Roy (2017) show non-vacuous and vacuous PAC-Bayes bounds for neural networks trained on structured and noisy MNIST, respectively. Smith & Le (2018) demonstrate benign overfitting for logistic regression on MNIST, interpreting the results using Bayesian Occam factors (MacKay, 2003). Several studies analyze two-layer networks (e.g., Cao et al., 2022; Kou et al., 2023). Wilson & Izmailov (2020) exactly reproduce the experiments in Zhang et al. (2016) with Gaussian processes and Bayesian neural networks, and explain the results using marginal likelihood. Bartlett et al. (2020) show that linear regression models can reproduce benign overfitting. They understand this phenomenon by studying the rank of the data covariance matrix, and minimum-norm least squares solutions, related to how Maddox et al. (2020) explain double descent through effective dimensionality. We will return to this reasoning in the next sections on overparametrization and double descent.
@@ 5 Overparametrization
We are not interested in the parameters in isolation, but rather how the parameters control the properties of the functions we use to fit the data. We have already seen how arbitrarily large polynomials do not overfit the data, as long as they have a simplicity bias. Gaussian processes also provide compelling examples. A GP with an RBF kernel can be derived from an infinite sum of densely dispersed radial basis functions $\phi_{i}$: $f(x,w)=\sum_{i=1}^{\infty}w_{i}\phi_{i}(x)$ (MacKay, 1998). Similarly, using central limit theorem arguments, we can derive GP kernels corresponding to infinite single and multi-layer neural networks (Neal, 1996b; Lee et al., 2017) (the first of these being an infamous NeurIPS rejection!). Indeed, GPs are typically more flexible than any standard neural network, but often have their strongest performance relative to other model classes on small datasets, due a strong (but soft) simplicity bias.
@@ 5.2 Why does increasing parameters help performance?
There is often a perceived tension between flexibility and inductive biases, with the assumption that more flexible models must have weaker inductive biases. But as we have seen, not only is there not necessarily a trade-off between flexibility and inductive biases, the larger and more flexible models often have stronger inductive biases, which we illustrate in Figure 6.
@@ 6 Double Descent
Double descent can also be understood. As the number of parameters grows, initially the ability to fit the data improves, and the learned parameters have higher effective dimensionality (Section 3.2). Once the number of parameters has increased to the point we are choosing between parameter settings that all achieve a perfect data fit, the value of the loss is no longer the deciding factor of which parameters are selected. As we continue to increase the number of parameters, the volume of flat solutions grows, making these solutions more discoverable during training. The effective dimensionality of the solutions will thus decrease, and generalization will improve.
This explanation applies to linear models and neural networks alike, but we can gain additional insight with linear models. Suppose we have $Xw=y$, where $X$ is an $n\times d$ matrix of features, $w$ represents $d$ parameters, and $y$ are the $n$ datapoints. Once $d>n$, the model can interpolate the data perfectly with infinitely many parameter settings $w$. As $d$ continues to increase, the number of undetermined parameters, and flat directions in the loss, increases. The least squares solution $w^{*}=(X^{\top}X)^{-1}X^{\top}y$. For $d>n$ there are at most $n$ non-zero eigenvalues of the Hessian $X^{\top}X$. As $d$ continues to increase, the signal will get spread out across more parameters, causing individual parameters to be less strongly determined and the effective dimensionality to decrease. Alternatively, $w^{*}$ provides the minimum $\ell_{2}$ norm solution, favouring simpler models that rely primarily on the most informative directions in feature space (Bartlett et al., 2020). As with benign overfitting, these notions of simplicity can be formally characterized using countable hypothesis or PAC-Bayes bounds.
Following Maddox et al. (2020), in Figure 1 (bottom) we show double descent for a ResNet-18 and a linear model. For the ResNet we show cross-entropy loss on CIFAR-100 as we increase the width of each layer. We show mean-squared error for the linear model, which uses the weakly informative features $y+\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$. Both follow a similar trend: effective dimensionality increases up to the interpolation regime, and then decreases, at which point generalization and effective dimensionality are aligned.
It is also possible to track double descent with formal PAC-Bayes bounds, as in Lotfi et al. (2022a, Figure 7). In the second descent, larger models achieve similar empirical risk, but can be more compressible.
@@ 8.2 Universal Learning
In other words, not only are neural networks learning rich representations of data, they are learning representations that are relatively universal across real-world problems, compared to other model classes. We emphasize that for in-context learning, these are not fixed representations. During pre-training, transformers learn to learn, discovering inductive principles such as Occam’s razor (Gruver et al., 2024; Goldblum et al., 2024). In the GP analogy, we can think of the pre-trained transformer as a large mixture of GP experts, with different kernels. Based on the downstream dataset, the transformer selects the appropriate combination of kernels, based on what it has seen in pre-training.
@@ 9 Discussion
What is the role of the optimizer in deep learning generalization? There is a conventional wisdom that the green and pink colours in Figure 6 are essentially inverted, and that the main reason deep learning works is because the implicit biases of stochastic optimizers cause them to traverse a relatively small subspace of low loss solutions with good generalization. However, it has been shown that not only full batch gradient descent, but even guess and check, stopping when the loss falls below a threshold, can find solutions with similar generalization as stochastic optimization (Geiping et al., 2021; Chiang et al., 2022), in alignment with Figure 6 (right). While in principle it is possible for an optimizer to still find bad optima under such a loss landscape, it would have to be actively adversarial. Far from adversarial, stochastic optimization has biases that can indeed improve generalization. But, importantly, these biases are not necessary for respectable generalization. Of course, stochastic optimization is much more computationally practical than the alternatives. No one is suggesting we use guess and check! Moreover, developing optimizers which generalize better under a given computational budget is a particularly exciting research direction, especially with recent results showing the rise of second-order optimizers (Liu et al., 2025; Vyas et al., 2024). Finally, the generalization bounds of Section 3.1 can be evaluated regardless of whether the model uses stochastic optimization, and indeed these bounds track the benign overfitting behaviour of Gaussian processes, which perform Bayesian inference.
@@
What is the relationship between structural risk minimization and soft inductive biases? SRM is a way to encode a soft inductive bias, but is more narrowly focused, and often differently motivated. SRM is often used as a mechanism to reduce VC dimension, trading off data fit with model complexity. It is not typically used as a prescription for arbitrarily flexible models, and indeed model selection tools with priors corresponding to standard $\ell_{2}$ regularization suggest we should use intermediate order models (Bishop, 2006). A key point in our paper is that we can embrace models that fit data perfectly (including noise) but still have a bias for simplicity. Other ways of implementing soft inductive biases include overparametrization, Bayesian priors and marginalization, the optimizer, and architectural specification.
@@
How can we better understand generalization in deep learning? There are many fascinating open questions in deep learning generalization. As an approach, we believe it is promising to analyze the solutions neural networks actually reach to explain their behaviours. The generalization bounds of Section 3.1 are fully empirical, non-asymptotic, and can be evaluated using a single sample. We view being able to empirically evaluate the bounds as essential in determining how much of the empirical model behaviour is actually explained by the theory. We have found the Solomonoff prior particularly useful for evaluating descriptive generalization bounds. Solomonoff induction uses a maximally overparametrized model, containing every possible program, but formalizes an ideal learning system that assigns exponentially higher weights to shorter programs. In the future, it would be enlightening to investigate properties of priors that may lead to tighter bounds, ever more closely describing deep learning generalization.

### v2

@@ Abstract
Deep neural networks are often seen as different from other model classes by defying conventional notions of generalization. Popular examples of anomalous generalization behaviour include benign overfitting, double descent, and the success of overparametrization. We argue that these phenomena are not distinct to neural networks, or particularly mysterious. Moreover, this generalization behaviour can be intuitively understood, and rigorously characterized, using long-standing generalization frameworks such as PAC-Bayes and countable hypothesis bounds. We present soft inductive biases as a key unifying principle in explaining these phenomena: rather than restricting the hypothesis space to avoid overfitting, embrace a flexible hypothesis space, with a soft preference for simpler solutions that are consistent with the data. This principle can be encoded in many model classes, and thus deep learning is not as mysterious or different from other model classes as it might seem. However, we also highlight how deep learning is relatively distinct in other ways, such as its ability for representation learning, phenomena such as mode connectivity, and its relative universality.
@@ 1 Introduction
We will also treat all of these phenomena collectively, through a unifying notion of soft inductive biases. While inductive biases are often thought of as restriction biases — constraining the size of a hypothesis space for improved data efficiency and generalization — there is no need for restriction biases. Instead, we can embrace an arbitrarily flexible hypothesis space, combined with soft biases that express a preference for certain solutions over others, without entirely ruling out any solution, as illustrated in Figure 3. Frameworks such as PAC-Bayes embody this view of inductive biases, capable of producing non-vacuous generalization bounds on models with even billions of parameters, as long as these models have a prior preference for certain solutions over others (Lotfi et al., 2024b). Broadly speaking, a large hypothesis space, combined with a preference for simple solutions, provides a provably useful recipe for good performance, as in Figure 2.
There are also other phenomena of recent interest, such as scaling laws and grokking, which are not our focus, because these are not typically treated as inconsistent with generalization theory, or distinct to neural networks. However, we note the PAC-Bayes and countable hypothesis generalization frameworks of Section 3 also describe LLMs, and even Chinchilla scaling laws (Hoffmann et al., 2022; Finzi et al., 2025). Moreover, deep learning of course is different in other ways. In Section 8, we discuss relatively distinctive features of deep neural networks, such as representation learning, mode connectivity, and broadly successful in-context learning.
@@ 2 Soft Inductive Biases
In Residual Pathway Priors (RPP) (Finzi et al., 2021), it was shown that a soft bias for equivariance constraints is often as effective as a model that had been perfectly constrained for a given problem. For example, a soft bias for rotation equivariance would work as well as a rotationally equivariant model for molecules, which are rotation invariant. After exposure to only a very small amount of data, the soft bias would converge to near-perfect rotation equivariance, since the model is encouraged (but not constrained) to represent the data with symmetries, and it can do so exactly, even with a small amount of data. Moreover, in cases where the data only contained an approximate symmetry, or no symmetry at all, the RPP approach would significantly outperform a model with hard symmetry constraints.
@@ 3 Generalization Frameworks
It turns out these phenomena are also formally characterized by generalization frameworks that have existed for many decades, including PAC-Bayes (McAllester, 1999; Guedj, 2019; Alquier et al., 2024) and simple countable hypothesis bounds (Valiant, 1984; Shalev-Shwartz & Ben-David, 2014). We introduce these frameworks in Section 3.1. We then define effective dimensionality in Section 3.2 which we will return to later in the paper for intuition. Finally, we introduce frameworks that do not describe these phenomena in Section 3.3, but have greatly impacted the conventional wisdom in thinking about generalization.
@@ 3.1 PAC-Bayes and countable hypothesis bounds
This bound is related to the finite hypothesis bound, but includes a prior $P(h)$ and a countable rather than finite hypothesis space (Ch 7.3, Shalev-Shwartz & Ben-David, 2014). We can think of the prior as a weighting function that weights certain hypotheses more highly than others. Importantly, we can use any prior to evaluate the bound: it need not have generated the true hypothesis for the data, contain the true hypothesis, or even be used by the model that is trained to find some hypothesis $h^{*}$. If the model uses a prior quite different from the prior used to evaluate Eq. (1), then the bound will simply become loose. We include an elementary proof of this bound in Appendix C.
We can derive informative bounds through a Solomonoff prior $P(h)=2^{-K(h|A)}/Z$ (Solomonoff, 1964), where $K$ is the prefix-free Kolmogorov complexity of $h$ taking as input model architecture $A$, and the normalizing constant $Z\leq 1$ by the Kraft inequality (Kraft, 1949). Substituting this prior into Eq. (1),
@@ 3.2 Effective Dimensionality
Solutions with lower effective dimensionality are flatter, meaning that the associated parameters can be perturbed without significantly increasing the loss. Flatness is not the only factor influencing generalization, and flatness as measured by the Hessian is not parametrization invariant meaning it is easy to find and construct examples where flatter solutions do not generalize better (e.g., Dinh et al., 2017). On the other hand, many standard procedures are not parametrization invariant (e.g., SGD and $\ell_{2}$ regularization), and the connection between flatness and generalization is not a spurious empirical association. We have a mechanistic understanding of why flatness can lead to better generalization: flatter solutions are more compressible, have better Occam factors, tend to lead to wider decision boundaries, and tighter generalization bounds (Hinton & Van Camp, 1993; Hochreiter & Schmidhuber, 1997; MacKay, 2003; Keskar et al., 2016; Izmailov et al., 2018; Foret et al., 2020; Maddox et al., 2020).
@@
We will return to effective dimensionality for intuition when discussing generalization phenomena.
@@ 3.3 Other Generalization Frameworks
Rademacher complexity (Bartlett & Mendelson, 2002) exactly measures the ability for a model to fit uniform $\{+1,-1\}$ random noise. Similarly, the VC dimension (Vapnik et al., 1994) measures the largest integer $d$ such that the hypothesis space $\mathcal{H}$ can fit (“shatter”) any set of $d$ points with $\{+1,-1\}$ labels. The fat-shattering dimension (Alon et al., 1997) $\text{fat}_{\gamma}(\mathcal{H})$ refines the VC dimension to fitting (“shattering”) labels by some margin $\gamma$. Unlike PAC-Bayes, all of these frameworks penalize the size of the overall hypothesis space $\mathcal{H}$, suggesting a prescription for restriction biases, rather than the soft inductive biases of Section 2. We discuss these frameworks further in Appendix B, with a comparative summary in Table 1. We note that while Rademacher complexity in itself does not describe these phenomena, it could be combined with a prior that would lead to informative bounds for expressive model classes.
@@ 4 Benign Overfitting
There is by now a large body of work studying and reproducing benign overfitting with other model classes. Yet the conventional wisdom of benign overfitting as a mysterious and deep learning specific phenomenon, one that still requires rethinking generalization, persists. It is not our intention, nor would it be possible, to cover all of this work here, but we note some of the key developments. Dziugaite & Roy (2017) show non-vacuous and vacuous PAC-Bayes bounds for neural networks trained on structured and noisy MNIST, respectively. Smith & Le (2018) demonstrate benign overfitting for logistic regression on MNIST, interpreting the results using Bayesian Occam factors (MacKay, 2003). Several studies analyze two-layer networks (e.g., Cao et al., 2022; Kou et al., 2023). Wilson & Izmailov (2020) exactly reproduce the experiments in Zhang et al. (2016) with Gaussian processes and Bayesian neural networks, and explain the results using marginal likelihood. Bartlett et al. (2020) show that linear regression models can reproduce benign overfitting. They understand this phenomenon by studying the rank of the data covariance matrix, and minimum-norm least squares solutions.
@@ 5 Overparametrization
We are not interested in the parameters in isolation, but rather how the parameters control the properties of the functions we use to fit the data. We have already seen how arbitrarily large polynomials do not overfit the data, as long as they have a simplicity bias. Gaussian processes also provide compelling examples. A GP with an RBF kernel can be derived from an infinite sum of densely dispersed radial basis functions $\phi_{i}$: $f(x,w)=\sum_{i=1}^{\infty}w_{i}\phi_{i}(x)$ (MacKay, 1998). Similarly, using central limit theorem arguments, we can derive GP kernels corresponding to infinite single and multi-layer neural networks (Neal, 1996b; Lee et al., 2017; Matthews et al., 2018) (the first of these being an infamous NeurIPS rejection!). Indeed, GPs are typically more flexible than any standard neural network, but often have their strongest performance relative to other model classes on small datasets, due a strong (but soft) simplicity bias.
@@ 5.2 Why does increasing parameters help performance?
There is often a perceived tension between flexibility and inductive biases, with the assumption that more flexible models must have weaker inductive biases. But as we have seen, the larger and more flexible models often have stronger inductive biases, which we illustrate in Figure 6.
@@ 6 Double Descent
Following Maddox et al. (2020), consider Figure 1 (bottom left), showing cross-entropy loss on CIFAR-100 with increases in the width of each layer of a ResNet-18, training to convergence. In the underparametrized regime (yellow), train and test loss both decrease, as increases in flexibility enable the model to capture more useful information content in the data, which increases the effective dimensionality (Section 3.2) of the Hessian of the trained parameters. In the the transition regime (pink), the training loss is still decreasing, and the information content in the trained parameters of the model is still increasing, leading to a continued increase in effective dimensionality. However, the way the model is capturing structure is leading to some overfitting, increasing the test loss. In the interpolation regime (green), where number of parameters exceed the number of data points, test loss again decreases with increases in parameters. Importantly, in this second descent, all models achieve a perfect fit to the data, and therefore increased flexibility cannot be the reason that the models with more parameters achieve better generalization. Instead, as we continue to increase the number of parameters, the volume of compressible flat solutions grows, making these solutions more discoverable during training. The effective dimensionality of the solutions thus decreases, and generalization will improve.
In Figure 1 (bottom right) we show double descent for the mean-squared error of a linear model, $Xw=y$, where $X$ is an $n\times d$ matrix of features, $w$ represents $d$ parameters, and $y$ are the $n$ datapoints. This model uses weakly informative features $y+\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$. Once $d>n$, the model can interpolate the data perfectly with infinitely many parameter settings $w$. Bartlett et al. (2020) shows that the least squares solution $w^{*}=(X^{\top}X)^{-1}X^{\top}y$ provides the minimum $\ell_{2}$ norm solution amongst parameter settings that perfectly fit the data, favouring simpler models that rely primarily on the most informative directions in the feature space. We can also consider the effective dimensionality of the parameter covariance matrix (inverse Hessian), or, equivalently, the number of relatively small eigenvalues of the Hessian $X^{\top}X$. Using the Marchenko-Pastur distribution in random matrix theory (Marchenko & Pastur, 1967; Dobriban & Wager, 2018), we can predict in this setting that the small eigenvalues will get larger as the number of parameters $d$ increases past $n$, which in turn decreases the variance error (in the bias-variance decomposition) of the model (Hastie et al., 2022), leading to better generalization. Curth et al. (2023) provides a detailed analysis of double descent in classical statistical models, including a generalized notion of effective parameters.
As with benign overfitting, these notions of simplicity can be formally characterized using countable hypothesis or PAC-Bayes bounds. It is also possible to track double descent with formal PAC-Bayes bounds, as in Lotfi et al. (2022a, Figure 7). In the second descent, larger models achieve similar empirical risk, but can be more compressible.
@@ 8.2 Universal Learning
In other words, not only are neural networks learning rich representations of data, they are learning representations that are relatively universal across real-world problems, compared to other model classes. We emphasize that for in-context learning, these are not fixed representations. During pre-training, transformers learn to learn, discovering inductive principles such as Occam’s razor (Gruver et al., 2024; Goldblum et al., 2024). In the GP analogy, we can think of the pre-trained transformer as a large mixture of GP experts, with different kernels. Conditioned on the downstream dataset, the transformer selects the appropriate combination of kernels, based on what it has seen in pre-training.
@@ 9 Discussion
What is the role of the optimizer in deep learning generalization? There is a conventional wisdom that the green and pink colours in Figure 6 are essentially inverted, and that the main reason deep learning works is because the implicit biases of stochastic optimizers cause them to traverse a relatively small subspace of low loss solutions with good generalization. However, it has been shown that full batch gradient descent, and even guess and check, stopping when the loss falls below a threshold, can find solutions with similar or only slightly worse generalization as stochastic optimization (Geiping et al., 2021; Chiang et al., 2022), in alignment with Figure 6 (right). While in principle it is possible for an optimizer to still find bad optima under such a loss landscape, it would have to be actively adversarial. Far from adversarial, stochastic optimization has biases that can indeed improve generalization. But, importantly, these biases are not necessary for respectable generalization. Of course, stochastic optimization is much more computationally practical than the alternatives. No one is suggesting we use guess and check! Moreover, developing optimizers which generalize better under a given computational budget is a particularly exciting research direction, especially with recent results showing the rise of second-order optimizers (Liu et al., 2025; Vyas et al., 2024). Finally, the generalization bounds of Section 3.1 can be evaluated regardless of whether the model uses stochastic optimization, and indeed these bounds track the benign overfitting behaviour of Gaussian processes, which perform Bayesian inference.
@@
What is the relationship between structural risk minimization and soft inductive biases? SRM is a way to encode a soft inductive bias, but is more narrowly focused, and often differently motivated. SRM is often used as a mechanism to reduce VC dimension, trading off data fit with model complexity. It is not typically used as a prescription for arbitrarily flexible models, and indeed model selection tools with priors corresponding to standard $\ell_{2}$ regularization suggest we should use intermediate order models (Bishop, 2006). A key point in this paper is that we can embrace models that fit data perfectly (including noise) but still have a bias for simplicity. Other ways of implementing soft inductive biases include overparametrization, Bayesian priors and marginalization, the optimizer, and architectural specification.
Should we use a different sized model depending on how much pre-training data are available? Assuming the dataset is fixed (but of arbitrary size), and we want the best performance, then no. Our beliefs about the process that generated the data are typically independent of how many training points we happen to have access to, and we should honestly represent our beliefs in our model. A large model can achieve good performance for all data sizes, as long as it combined with a simplicity bias, as we saw in the polynomial example.
Can restriction biases have computational benefits? While we are primarily considering the principles of model construction in an idealized setting, where computation is not a constraint, restriction biases such as sparsity and parameter sharing can be a practical design decision for computational reasons. However, even when considering computation, restriction biases can be undesirable. Indeed, recent work shows that parameter sharing can be a poor principle for compute-optimal scaling (Potapczynski et al., 2024).
@@
How can we better understand generalization in deep learning? There are many fascinating open questions in deep learning generalization. As an approach, we believe it is promising to analyze the solutions neural networks actually reach to explain their behaviours. The generalization bounds of Section 3.1 are fully empirical, non-asymptotic, and can be evaluated using a single sample. We view being able to empirically evaluate the bounds as essential in determining how much of the empirical model behaviour is actually explained by the theory. We have found the Solomonoff prior particularly useful for evaluating descriptive generalization bounds. Solomonoff induction uses a maximally overparametrized model, containing every possible program, but formalizes an ideal learning system that assigns exponentially higher weights to shorter programs. In the future, it would be enlightening to investigate properties of priors that may lead to tighter bounds, ever more closely describing deep learning generalization. It would be particularly exciting to move beyond Kolmogorov complexity as a measure of information content, in order to distinguish between incompressibility due to randomness versus incompressibility due to structural complexity.